OpenStudy (anonymous):

square root of 2x-3=7 find x

OpenStudy (anonymous):

Square both sides

OpenStudy (anonymous):

i dont know how to do it at all

OpenStudy (anonymous):

First you add 3 to the 7 and that equals what?

OpenStudy (anonymous):

After you find the answer your equation should look like this 2x = ___

OpenStudy (anonymous):

Please clarify, is it \(\sqrt{2x-3}=7\) or \(\sqrt{2x}-3=7\)?

OpenStudy (pfenn1):

Is the equation\[\sqrt{2}x-3=7\]or\[\sqrt{2x}-3=7\]or\[\sqrt{2x-3}=7\]?

OpenStudy (anonymous):

Then you divide both sides by 2. So 2x/2 = (7+3)/2.

OpenStudy (anonymous):

Now you have x. Then you take the square root of x. Tadaaa!!! :)

OpenStudy (anonymous):

so 5?

OpenStudy (anonymous):

Yes, but what is the square root of 5? Thats your answer. Which equation above is it? Look at pfenn's question above my answer. Which one does your equation look the most like?